# A Survey on Recommendation Unlearning: Fundamentals, Taxonomy, Evaluation, and Open Questions

### v1

@@ III-A1 Model-agnostic Methods
Current reverse unlearning methods in recommender systems mainly rely on influence function [46, 47, 7] to estimate the influence of target data on model parameters. Specifically, it estimates the influence by weighting a data record by $\epsilon$. Formally, the $\epsilon$-weighted model parameter is
@@
According to [46], leverage second-order Talor expansion, the estimated influence of $z$ is given by
@@
where $\nabla_{\theta}\ell(z,\theta_{0})$ is the gradient vector, and $H_{\theta_{0}}:=\sum_{i=1}^{n}\nabla^{2}_{\theta}\ell(z,\theta_{0})$ is the Hessian matrix and is positive definite by assumption. The derivation in Eq.(6) is equivalent to a single step of the Newton optimization update. Therefore, influence function-based reverse unlearning methods can be interpreted as performing a one-step reverse Newton update. Zhang et al. [121] directly apply this closed-form one-step reverse Newton update to the matrix factorization model. There are also other tasks in recommender systems that utilize the influence function. For example, Wang et al. [105] study the user-controllable recommendation task, where users can control which interactions are used for training. This can also be formulated as a recommendation unlearning task.
@@
Firstly, due to the large number of users and items in recommendation tasks, the user and item embeddings form high-dimensional dense matrices, which incur considerable computational overhead when explicitly calculating the influence function. As shown in Eq.(6), this calculation involves computing the inverse of the Hessian matrix and then multiplying it by the gradient vector, which has a computational complexity of $O(n^{2})$. Since the forget set typically accounts for a small portion of the training data, Li et al. [58] found that the parameter changes before and after unlearning primarily affect the embeddings of the unlearning target. Changes to other parameters (e.g., non-targeted embeddings and parameters unrelated to embeddings) are negligible. Consequently, they propose selectively calculating the influence function for the target embeddings only, thereby reducing computational overhead at the root level. Similarly, Zhang et al. [122] propose an importance-based pruning strategy to reduce the model size, thereby lowering the overall computational overhead of the influence function. The importance is estimated based on the neighborhood relationships among the unlearning targets.
Secondly, influence function is essentially an approximation, and its accuracy is not always guaranteed, particularly for deep learning models [6]. In recommendation tasks, removing the influence of target data can have a cascading effect on neighboring data, thereby further compromising model utility. To address this issue, Li et al. [58] incorporate a collaborative term into Eq. (3) to mitigate this negative effect. Instead of directly removing the target data, this collaborative term replaces it with the average rating of the neighboring data, which compensates for the inaccurate estimation of influence function.
@@
  • The derivation of the influence function assumes that only a small portion of the training data is removed. The tolerance for the volume of the forget set is not comparable to exact unlearning. Moreover, in real-world scenarios where unlearning requests are submitted sequentially and must be processed immediately (i.e., sequential/stream unlearning [29]), the assumption that only a small portion of the data is unlearned no longer holds.
@@
Lie et al. [66] first propose active unlearning in recommendation tasks by fine-tuning the original model on the retain set, which is faster than retraining the model from scratch. Furthermore, they apply a 2nd-order gradient optimization technique to further enhance the efficiency of this unlearning process.
Alshehri et al. [3] propose active unlearning by flipping the labels of data in forget set. However, fine-tuning the model with label-flipped records may unintentionally overwrite the parameters of the original model, negatively impacting model utility for other users. To address this issue, they also sample data from the retain set during fine-tuning.
Label flipping is limited to binary ratings $\{0,1\}$, which presents challenges when dealing with value-based ratings. To overcome this issue, Sinha et al. [94] propose flipping the loss function instead of the labels, reversing the direction of the loss (i.e., changing addition to subtraction). They also utilize data from the retain set to prevent over-unlearning. These two loss functions (i.e., flipped loss with the forget set and original loss with the retain set) are linearly combined with a balancing coefficient. Similarly, You et al. [118] also flip the loss function and tune it on the forget set. The key difference from previous methods is that they solely rely on the flipped loss, updating it in a single step using the Fisher information matrix. This approach does not depend on the retain set, making it more convenient and efficient for implementation in real-world applications.  Dang et al. fine-tune the slipped loss by introducing an additional module to enhance utility [21].
In addition to fine-tuning, Chaturvedi et al. [14] propose a model fusion approach to achieve active unlearning. Specifically, they fine-tune one model using noisy data and another, smaller model using the retained data set. They then apply a convolutional fusion function to combine the models, thereby efficiently enabling unlearning.
@@
Another emerging direction for achieving unlearning also takes a functional perspective, but with an approach distinct from “learning to unlearn”. Liu et al. [65] reveal that collaborative filtering can be formulated as a mapping-based approach, where the recommendations are obtained by multiplying the user-item interaction matrix with a mapping matrix. This formulation simplifies unlearning to the manipulation of the mapping matrix. While this method provides valuable insights into model-agnostic recommendation unlearning, it has some limitations. The arbitrary approximation of recommendation models as mapping-based approaches lacks a solid theoretical foundation, making it highly dependent on the accuracy of the mapping matrix approximation.
@@ III-A2 Model-specific Methods
  • Bi-linear recommendation model: Xu et al. [115] propose an exact unlearning method for bilinear recommendation models, which utilize alternating least squares for optimization [35]. The core idea of their approach is fine-tuning. In these models, the confidence matrix, which is multiplied by the predicted ratings, plays a key role. By setting the target elements of this matrix to zero during fine-tuning, the method effectively performs exact unlearning. However, the authors also note that exact unlearning may not always be feasible in real-world applications. This is because additional optimization techniques, such as early stopping, are often employed, which can introduce complexities that prevent exact unlearning from being fully realized.
  • KNN-based recommendation model: $k$-Nearest Neighbor (KNN) is widely used in a variety of recommendation scenarios [85, 74, 4] due to its several key advantages. These include its transparency and explainability, cost-effectiveness in scaling to industrial workloads, and significantly lower training time. KNN’s simplicity and efficiency make it an attractive choice for both research and practical applications in the field of recommender systems. Unlike the models discussed in the review, KNN is a non-parametric model, which inherently facilitates completeness by simply removing the unlearning target. Schelter et al. [89] propose an efficient indexing method to accelerate this unlearning process, making it both faster and more scalable.
  • Graph-based recommendation model: Graph Neural Network (GNN)-based recommendation models have gained prominence in recent years [33, 81, 125]. Hao et al. [30] propose inserting a learnable delete operator at each layer of the GNN and fine-tuning the model with a linear combination of two objectives. The first component focuses on unlearning the target data, while the second component aims to maintain consistent feature representation, thereby preserving the model’s utility.
  • Knowledge graph-based recommendation model: Knowledge graph-based recommendation models utilize domain-specific knowledge to produce recommendations, making them a crucial category of systems, especially for tackling the cold-start problem. These models apply rules, reasoning, and constraints derived from domain knowledge [50]. In a knowledge graph, a record is represented as a triple $<s,p,o>$ where $s$, $p$, and $o$ denote the subject, predicate, and object, respectively. Wang et al. [107] propose two types of unlearning for knowledge graph-based recommendation models: (i) passive forgetting, which unlearns the target data based on user requests, and (ii) intentional forgetting, which optimizes the entire knowledge graph. Their unlearning is achieved through rule replacement.
@@ III-A3 Scenario-specific Methods
  • Federated Recommendation: The federated learning paradigm enables clients to collaboratively train a recommendation model without sharing their original data. Although federated learning offers a safeguard for user privacy by preventing data sharing, it cannot achieve unlearning. In federated recommendation models, each user typically acts as an individual client. Yuan et al. [119] propose an unlearning method tailored for federated recommender systems. Building on the design of FedEraser [63], their approach efficiently recalculates the update gradients for the retain clients by conducting a few-step fine-tuning process to determine gradient direction, which is then scaled by the original update length.
@@
  • Sequential Recommendation: Sequential recommender systems, which leverage clients’ sequential product browsing history, have become indispensable for providing personalized product recommendations. These systems track the order in which users interact with products, allowing them to predict future preferences and suggest relevant items [82]. Ye et al. [117] build upon the idea of active unlearning by introducing noise into the interaction sequences. They then fine-tune the original model using these noisy sequences to effectively achieve unlearning.
@@
  • Session-based Recommendation: Session-based recommendation models have proven effective in predicting users’ future interests by leveraging their recent sequential interactions. While these models also emphasize sequence-aware data, the key distinction of session-based approaches lies in their exclusive focus on the current user session [112]. Typically, these methods do not rely on long-term user information, tailoring recommendations solely based on the immediate session context. Xin et al. [113] focus on unlearning an item from the current session. They follow the idea of exact unlearning and implement an ensemble retraining framework. Similar to the approach in [15], they use balanced clustering for session division and apply an attention network for aggregating sub-models.
@@
  • LLM-based Recommendation: LLMs, with their vast number of parameters and training on extensive datasets, exhibit remarkable performance across a wide range of tasks [80, 101, 23]. The impressive advancements in LLMs have also sparked interest in using them as recommender systems. The effectiveness of LLM-based recommender systems lies in the models’ extensive open-world knowledge and their reasoning capabilities [2, 61]. LLM-based recommender systems acquire recommendation capabilities through instruction tuning on user interaction data. Note that due to the high cost of retraining LLMs, the efficiency of LLM-based recommendation unlearning has become a significant concern. To address this issue, Wang et al. [104] leverage Low-Rank Adaptation (LoRA) for lightweight fine-tuning. Building on the idea of active unlearning, and following a similar approach to [20], they adopt a teacher-student framework. In this framework, a remembering teacher preserves knowledge from the retain set, while a forgetting teacher is used to unlearn information from the forget set. Knowledge distillation is then performed using a Kullback-Leibler divergence loss. Hu et al.[39] also leverage LoRA for fine-tuning, but they follow the idea of exact unlearning (i.e., ensemble retraining framework). Rather than training separate sub-models, which is computationally prohibitive for LLM-based recommendation models, they train an adapter (LoRA module) for each data subset. This approach enables exact unlearning in LLM-based recommender systems, allowing for highly efficiency retraining.
@@ III-B1 In-training Unlearning
Attribute unlearning was first introduced by [26], focusing on unlearning attributes in MultVAE. The core idea is adversarial training, which can be extended to other recommendation models as well. Specifically, they incorporated an adversarial decoder into MultVAE, which acts as an attacker attempting to infer user attributes. The encoder, in turn, works to deceive the adversarial decoder through bi-level optimization. As a result, the model learns embeddings that are robust to attribute inference attacks while still effectively modeling user preferences.
Escobedo et al. [24] propose an alternative approach to in-training attribute unlearning through stereotypicality-based obfuscation. Specifically, they manipulate interaction records by imputation, removal, and weighting, all guided by a stereotypicality score. Interaction records with a high stereotypicality score indicate a strong association with certain attribute information. For example, a male user who predominantly listens to music tracks associated with male stereotypes. By obfuscating these records, they aim to eliminate attribute information while minimizing any negative impact on model utility.
@@ III-B2 Post-training Unlearning
Li et al. [59] propose a bi-objective optimization framework for achieving post-training attribute unlearning. The first objective (i.e., distinguishability loss) is directly related to the primary goal of attribute unlearning. They design two types of distinguishability losses: user loss and distributional loss. The user loss manipulates the user embeddings such that users with the same attribute label are pushed apart, while users with different attribute labels are pulled together. The distributional loss treats users with the same attribute label as a group and minimizes the distance between their distributions. The second objective (i.e., regularization loss) supports the secondary goal of attribute unlearning. Since the training data is not available during unlearning, they introduce a regularization term between the original and unlearned models to help preserve model utility. The two objectives are combined through a linear combination, with a balancing coefficient to control the trade-off between them.
Their subsequent work further refines both objectives [18]. For the distributional loss, they introduce an anchor distribution during computation, which significantly enhances the efficiency when calculating multi-class attributes. Previously, the complexity of the distributional loss grows exponentially with the number of classes; in contrast, the anchor-based distributional loss scales linearly, leading to more efficient computation. For the regularization loss, prior work assumed that similarity in model parameters directly corresponds to alignment in model performance. However, this assumption does not always hold [8]. To address this issue, they introduce a regularization term in the functional space, which more accurately aligns the recommendation performance with that of the original model, while still avoiding access to the training data.
@@
  • A pluggable design is highly favored in industrial applications because it allows for more flexibility and modularity. Such a design enables the unlearning operation to be easily integrated or detached from the recommender system. Although post-training unlearning offers flexibility for unlearning execution, there is also a growing need for the unlearning process to be reversible. This is crucial, particularly when user preferences change over time or when new regulations or legal requirements emerge.
Takeaways. To enhance readability, we summarize the focused problems and the main techniques employed in each of the methods discussed above in Table I. This table provides a concise overview, allowing a quick grasp of the core issues and approaches utilized across the various methods, offering a valuable reference for understanding the landscape of research in the area of recommendation unlearning.
@@
Unlearning target Venue Focused problem Main techniques Model/Scenario
User/Item/Sample CSCWD’24 [55] Completeness Ensemble retraining (SISA) Agnostic
WWW’22 [15] Completeness Ensemble retraining Agnostic
NeurIPS’23 [56] Trade-off between efficiency and utility Ensemble retraining, optimal transport Agnostic
CIKM’23 [121] Efficiency Influence function Agnostic
ESWA’23 [58] Efficiency, utility Influence function Agnostic
TORS’24 [122] Efficiency Influence function, pruning Agnostic
AAAI’24 [105] Utility Influence function Agnostic
KBS’24 [57] Utility Ensemble retraining, curriculum learning Agnostic
AAAI’24 [118] Utility Loss flipping, Fisher information matrix Agnostic
TOIS’24 [21] Utility Loss flipping, multi-objective optimization Agnostic
Arxiv’22 [66] Efficiency 2nd-order fine-tuning Agnostic
Arxiv’23 [65] Efficiency Matrix correction Agnostic
OpenReview’24 [14] Efficiency Convolutional fusion function Agnostic
CIKM’24 [30] Utility Fine-tuning, bi-objective optimization Graph
Arxiv’23 [115] Completeness, efficiency Alternating least squares, fine-tuning Bi-linear
Arxiv’24 [94] Utility, multi-modal Loss flipping Multi-modal
Arxiv’24 [39] Efficiency, completeness LoRA, ensemble retraining LLM
FCS’25 [104] Efficiency, utility LoRA, teacher-student framework LLM
User WSDM’23 [119] Federated learning Update calibration Federated
SIGIR’23 [89] Efficiency Fast indexing KNN-based
AJCAI’23 [117] Utility Noise injection Sequential
BigData’23 [3] Utility, efficiency Label flipping Agnostic
Item WSDM’24 [113] Completeness Ensemble retraining Session-based
DATA’24 [107] Efficiency, utility Rule replacement Knowledge graph
Attribute SIGIR’22 [26] In-training Adversarial training VAE
ECML PKDD’24 [24] In-training Item obfuscation Agnostic
MM’23 [59] Post-training Bi-objective optimization Agnostic
TOIS’24 [18] Post-training, multi-class Bi-objective optimization Agnostic

### v2

@@ III-A1 Model-agnostic Methods
Current reverse unlearning methods in recommender systems mainly rely on influence function [koh2017understanding, koh2019accuracy, basu2020second] to estimate the influence of target data on model parameters. Specifically, it estimates the influence by weighting a data record by $\epsilon$. Formally, the $\epsilon$-weighted model parameter is
@@
According to [koh2017understanding], leverage second-order Taylor expansion, the estimated influence of $z$ is given by
@@
where $\nabla_{\theta}\ell(z,\theta_{0})$ is the gradient vector, and $H_{\theta_{0}}:=\sum_{i=1}^{n}\nabla^{2}_{\theta}\ell(z,\theta_{0})$ is the Hessian matrix and is positive definite by assumption. The derivation in Eq.(6) is equivalent to a single step of the Newton optimization update. Therefore, influence function-based reverse unlearning methods can be interpreted as performing a one-step reverse Newton update. Zhang et al. [zhang2023closed] directly apply this closed-form one-step reverse Newton update to the matrix factorization model. There are also other tasks in recommender systems that utilize the influence function. For example, Wang et al. [wang2024would] study the user-controllable recommendation task, where users can control which interactions are used for training. This can also be formulated as a recommendation unlearning task.
@@
Firstly, due to the large number of users and items in recommendation tasks, the user and item embeddings form high-dimensional dense matrices, which incur considerable computational overhead when explicitly calculating the influence function. As shown in Eq.(6), this calculation involves computing the inverse of the Hessian matrix and then multiplying it by the gradient vector, which has a computational complexity of $O(n^{2})$. Since the forget set typically accounts for a small portion of the training data, Li et al. [li2023selective] found that the parameter changes before and after unlearning primarily affect the embeddings of the unlearning target. Changes to other parameters (e.g., non-targeted embeddings and parameters unrelated to embeddings) are negligible. Consequently, they propose selectively calculating the influence function for the target embeddings only, thereby reducing computational overhead at the root level. Similarly, Zhang et al. [zhang2023recommendation] propose an importance-based pruning strategy to reduce the model size, thereby lowering the overall computational overhead of the influence function. The importance is estimated based on the neighborhood relationships among the unlearning targets.
Secondly, influence function is essentially an approximation, and its accuracy is not always guaranteed, particularly for deep learning models [basu2021influence]. In recommendation tasks, removing the influence of target data can have a cascading effect on neighboring data, thereby further compromising model utility. To address this issue, Li et al. [li2023selective] incorporate a collaborative term into Eq. (3) to mitigate this negative effect. Instead of directly removing the target data, this collaborative term replaces it with the average rating of the neighboring data, which compensates for the inaccurate estimation of influence function.
@@
  • The derivation of the influence function assumes that only a small portion of the training data is removed. The tolerance for the volume of the forget set is not comparable to exact unlearning. Moreover, in real-world scenarios where unlearning requests are submitted sequentially and must be processed immediately (i.e., sequential/stream unlearning [gupta2021adaptive]), the assumption that only a small portion of the data is unlearned no longer holds.
@@
Lie et al. [liu2022forgetting] first propose active unlearning in recommendation tasks by fine-tuning the original model on the retain set, which is faster than retraining the model from scratch. Furthermore, they apply a 2nd-order gradient optimization technique to further enhance the efficiency of this unlearning process.
Alshehri et al. [alshehri2023forgetting] propose active unlearning by flipping the labels of data in forget set. However, fine-tuning the model with label-flipped records may unintentionally overwrite the parameters of the original model, negatively impacting model utility for other users. To address this issue, they also sample data from the retain set during fine-tuning.
Label flipping is limited to binary ratings $\{0,1\}$, which presents challenges when dealing with value-based ratings. To overcome this issue, Sinha et al. [sinha2024multi] propose flipping the loss function instead of the labels, reversing the direction of the loss (i.e., changing addition to subtraction). They also utilize data from the retain set to prevent over-unlearning. These two loss functions (i.e., flipped loss with the forget set and original loss with the retain set) are linearly combined with a balancing coefficient. Similarly, You et al. [you2024rrl] also flip the loss function and tune it on the forget set. The key difference from previous methods is that they solely rely on the flipped loss, updating it in a single step using the Fisher information matrix. This approach does not depend on the retain set, making it more convenient and efficient for implementation in real-world applications. Dang et al. [dangefficient] fine-tune the slipped loss by introducing an additional module to enhance utility.
In addition to fine-tuning, Chaturvedi et al. [chaturvedi2024unlearning] propose a model fusion approach to achieve active unlearning. Specifically, they fine-tune one model using noisy data and another, smaller model using the retained data set. They then apply a convolutional fusion function to combine the models, thereby efficiently enabling unlearning.
@@
Another emerging direction for achieving unlearning also takes a functional perspective, but with an approach distinct from “learning to unlearn”. Liu et al. [liu2023recommendation] reveal that collaborative filtering can be formulated as a mapping-based approach, where the recommendations are obtained by multiplying the user-item interaction matrix with a mapping matrix. This formulation simplifies unlearning to the manipulation of the mapping matrix. While this method provides valuable insights into model-agnostic recommendation unlearning, it has some limitations. The arbitrary approximation of recommendation models as mapping-based approaches lacks a solid theoretical foundation, making it highly dependent on the accuracy of the mapping matrix approximation.
@@ III-A2 Model-specific Methods
  • Bi-linear recommendation model: Xu et al. [xu2023netflix] propose an exact unlearning method for bilinear recommendation models, which utilize alternating least squares for optimization [he2016fast]. The core idea of their approach is fine-tuning. In these models, the confidence matrix, which is multiplied by the predicted ratings, plays a key role. By setting the target elements of this matrix to zero during fine-tuning, the method effectively performs exact unlearning. However, the authors also note that exact unlearning may not always be feasible in real-world applications. This is because additional optimization techniques, such as early stopping, are often employed, which can introduce complexities that prevent exact unlearning from being fully realized.
  • KNN-based recommendation model: $k$-Nearest Neighbor (KNN) is widely used in a variety of recommendation scenarios [resnick1994grouplens, ludewig2021empirical, ariannezhad2023personalized] due to its several key advantages. These include its transparency and explainability, cost-effectiveness in scaling to industrial workloads, and significantly lower training time. KNN’s simplicity and efficiency make it an attractive choice for both research and practical applications in the field of recommender systems. Unlike the models discussed in the review, KNN is a non-parametric model, which inherently facilitates completeness by simply removing the unlearning target. Schelter et al. [schelter2023forget] propose an efficient indexing method to accelerate this unlearning process, making it both faster and more scalable.
  • Graph-based recommendation model: Graph Neural Network (GNN)-based recommendation models have gained prominence in recent years [he2020lightgcn, peng2022svd, zheng2021dgcn]. Hao et al. [hao2024general] propose inserting a learnable delete operator at each layer of the GNN and fine-tuning the model with a linear combination of two objectives. The first component focuses on unlearning the target data, while the second component aims to maintain consistent feature representation, thereby preserving the model’s utility.
  • Knowledge graph-based recommendation model: Knowledge graph-based recommendation models utilize domain-specific knowledge to produce recommendations, making them a crucial category of systems, especially for tackling the cold-start problem. These models apply rules, reasoning, and constraints derived from domain knowledge [le2023constraint]. In a knowledge graph, a record is represented as a triple <s,p,o><s,p,o> where $s$, $p$, and $o$ denote the subject, predicate, and object, respectively. Wang et al. [wang2024forgetting] propose two types of unlearning for knowledge graph-based recommendation models: (i) passive forgetting, which unlearns the target data based on user requests, and (ii) intentional forgetting, which optimizes the entire knowledge graph. Their unlearning is achieved through rule replacement.
@@ III-A3 Scenario-specific Methods
  • Federated Recommendation: The federated learning paradigm enables clients to collaboratively train a recommendation model without sharing their original data. Although federated learning offers a safeguard for user privacy by preventing data sharing, it cannot achieve unlearning. In federated recommendation models, each user typically acts as an individual client. Yuan et al. [yuan2023federated] propose an unlearning method tailored for federated recommender systems. Building on the design of FedEraser [liu2021federaser], their approach efficiently recalculates the update gradients for the retain clients by conducting a few-step fine-tuning process to determine gradient direction, which is then scaled by the original update length.
@@
  • Sequential Recommendation: Sequential recommender systems, which leverage clients’ sequential product browsing history, have become indispensable for providing personalized product recommendations. These systems track the order in which users interact with products, allowing them to predict future preferences and suggest relevant items [quadrana2018sequence]. Ye et al. [ye2023sequence] build upon the idea of active unlearning by introducing noise into the interaction sequences. They then fine-tune the original model using these noisy sequences to effectively achieve unlearning. Schelter et al. [schelter2024snarcase] propose a low-latency unlearning method for next-basket recommendation via incremental view maintenance.
@@
  • Session-based Recommendation: Session-based recommendation models have proven effective in predicting users’ future interests by leveraging their recent sequential interactions. While these models also emphasize sequence-aware data, the key distinction of session-based approaches lies in their exclusive focus on the current user session [wu2019session]. Typically, these methods do not rely on long-term user information, tailoring recommendations solely based on the immediate session context. Xin et al. [xin2024effectiveness] focus on unlearning an item from the current session. They follow the idea of exact unlearning and implement an ensemble retraining framework. Similar to the approach in [chen2022recommendation], they use balanced clustering for session division and apply an attention network for aggregating sub-models.
@@
  • LLM-based Recommendation: LLMs, with their vast number of parameters and training on extensive datasets, exhibit remarkable performance across a wide range of tasks [openai2023gpt, touvron2023llama, dubey2024llama]. The impressive advancements in LLMs have also sparked interest in using them as recommender systems. The effectiveness of LLM-based recommender systems lies in the models’ extensive open-world knowledge and their reasoning capabilities [acharya2023llm, lin2024data]. LLM-based recommender systems acquire recommendation capabilities through instruction tuning on user interaction data. Note that due to the high cost of retraining LLMs, the efficiency of LLM-based recommendation unlearning has become a significant concern. To address this issue, Wang et al. [wang2025towards] leverage Low-Rank Adaptation (LoRA) for lightweight fine-tuning. Building on the idea of active unlearning, and following a similar approach to [chundawat2023can], they adopt a teacher-student framework. In this framework, a remembering teacher preserves knowledge from the retain set, while a forgetting teacher is used to unlearn information from the forget set. Knowledge distillation is then performed using a Kullback-Leibler divergence loss. Hu et al.[hu2024exact] also leverage LoRA for fine-tuning, but they follow the idea of exact unlearning (i.e., ensemble retraining framework). Rather than training separate sub-models, which is computationally prohibitive for LLM-based recommendation models, they train an adapter (LoRA module) for each data subset. This approach enables exact unlearning in LLM-based recommender systems, allowing for highly efficiency retraining.
@@ III-B1 In-training Unlearning
Attribute unlearning was first introduced by [ganhor2022unlearning], focusing on unlearning attributes in MultVAE. The core idea is adversarial training, which can be extended to other recommendation models as well. Specifically, they incorporated an adversarial decoder into MultVAE, which acts as an attacker attempting to infer user attributes. The encoder, in turn, works to deceive the adversarial decoder through bi-level optimization. As a result, the model learns embeddings that are robust to attribute inference attacks while still effectively modeling user preferences.
Escobedo et al. [escobedo2024making] propose an alternative approach to in-training attribute unlearning through stereotypicality-based obfuscation. Specifically, they manipulate interaction records by imputation, removal, and weighting, all guided by a stereotypicality score. Interaction records with a high stereotypicality score indicate a strong association with certain attribute information. For example, a male user who predominantly listens to music tracks associated with male stereotypes. By obfuscating these records, they aim to eliminate attribute information while minimizing any negative impact on model utility.
@@ III-B2 Post-training Unlearning
Li et al. [li2023making] propose a bi-objective optimization framework for achieving post-training attribute unlearning. The first objective (i.e., distinguishability loss) is directly related to the primary goal of attribute unlearning. They design two types of distinguishability losses: user loss and distributional loss. The user loss manipulates the user embeddings such that users with the same attribute label are pushed apart, while users with different attribute labels are pulled together. The distributional loss treats users with the same attribute label as a group and minimizes the distance between their distributions. The second objective (i.e., regularization loss) supports the secondary goal of attribute unlearning. Since the training data is not available during unlearning, they introduce a regularization term between the original and unlearned models to help preserve model utility. The two objectives are combined through a linear combination, with a balancing coefficient to control the trade-off between them.
Their subsequent work further refines both objectives [chen2024post]. For the distributional loss, they introduce an anchor distribution during computation, which reduces computational complexity from exponential to linear growth with the number of classes. For the regularization loss, they address the parameter-performance misalignment issue [benjamin2018measuring] by introducing a functional-space regularization term that better preserves original model performance without requiring training data.
Although post-training unlearning offers flexibility for unlearning execution, there is also a growing need for the unlearning process to be reversible. This is crucial, particularly when user preferences change over time or when new regulations or legal requirements emerge. Feng et al. [feng2025plug] propose a pluggable erasure module to achieve attribute unlearning. Specifically, they optimize through rate distortion to develop a pluggable erasure module capable of efficiently performing reversible unlearning operations, alongside attribute unlearning. Yu et al. [yu2025lego] propose a pluggable framework that enables dynamic unlearning of multiple attributes through a two-stage process, i.e., initial embedding calibration followed by attribute combination. A pluggable design is highly favored in industrial applications because it allows for more flexibility and modularity.
@@
  • Training data is fundamental to building a model with stratified recommendation performance. However, post-training attribute unlearning cannot directly access this data. Existing methods primarily rely on regularization techniques applied to model parameters or recommendation rank lists. While these approaches attempt to address the absence of training data, they often lead to reduced recommendation accuracy. These limitations underscore the need for more effective strategies that enable attribute unlearning without compromising model performance.
Takeaways. To enhance readability, we summarize the focused problems, the main techniques employed, and code reproducibility (with “Yes” indicating the accessible source code link) of each method discussed above in Table I. This table provides a concise overview, allowing a quick grasp of the core issues and approaches utilized across the various methods, offering a valuable reference for understanding the landscape of recommendation unlearning.
@@
Unlearning target Venue Focused problem Main techniques Model/Scenario Code
User/Item/Sample CSCWD’24 [li2024enhancing] Completeness Ensemble retraining (SISA) Agnostic No
WWW’22 [chen2022recommendation] Completeness Ensemble retraining Agnostic Yes
NeurIPS’23 [li2023ultrare] Efficiency, utility Ensemble retraining, optimal transport Agnostic Yes
CIKM’23 [zhang2023closed] Efficiency Influence function Agnostic No
ESWA’23 [li2023selective] Efficiency, utility Influence function Agnostic No
TORS’24 [zhang2023recommendation] Efficiency Influence function, pruning Agnostic Yes
AAAI’24 [wang2024would] Utility Influence function Agnostic Yes
KBS’24 [li2024making] Utility Ensemble retraining, curriculum learning Agnostic No
AAAI’24 [you2024rrl] Utility Loss flipping, Fisher information matrix Agnostic No
TOIS’24 [dangefficient] Utility Loss flipping, multi-objective optimization Agnostic No
Arxiv’22 [liu2022forgetting] Efficiency 2nd-order fine-tuning Agnostic No
Arxiv’23 [liu2023recommendation] Efficiency Matrix correction Agnostic No
OpenReview’24 [chaturvedi2024unlearning] Efficiency Convolutional fusion function Agnostic No
CIKM’24 [hao2024general] Utility Fine-tuning, bi-objective optimization Graph Yes
Arxiv’23 [xu2023netflix] Completeness, efficiency Alternating least squares, fine-tuning Bi-linear No
Arxiv’24 [sinha2024multi] Utility, multi-modal Loss flipping Multi-modal Yes
Arxiv’24 [hu2024exact] Efficiency, completeness LoRA, ensemble retraining LLM No
FCS’25 [wang2025towards] Efficiency, utility LoRA, teacher-student framework LLM Yes
User WSDM’23 [yuan2023federated] Federated learning Update calibration Federated No
SIGIR’23 [schelter2023forget] Efficiency Fast indexing KNN-based Yes
AJCAI’23 [ye2023sequence] Utility Noise injection Sequential No
BigData’23 [alshehri2023forgetting] Utility, efficiency Label flipping Agnostic No
VLDB’24 [schelter2024snarcase] Efficiency Incremental view maintenance Sequential Yes
Item WSDM’24 [xin2024effectiveness] Completeness Ensemble retraining Session-based Yes
DATA’24 [wang2024forgetting] Efficiency, utility Rule replacement Knowledge graph No
Attribute SIGIR’22 [ganhor2022unlearning] In-training Adversarial training VAE Yes
BFIR’24 [escobedo2024simultaneous] In-training, multi-attribute Adversarial training VAE Yes
ECML PKDD’24 [escobedo2024making] In-training Item obfuscation Agnostic No
ACM MM’23 [li2023making] Post-training Bi-objective optimization Agnostic Yes
TOIS’24 [chen2024post] Post-training, multi-class Bi-objective optimization Agnostic No
WWW’25 [feng2025plug] Post-training, pluggable Rate distortion Agnostic No
ACM MM’25 [yu2025lego] Post-training, multi-attribute Calibration and combination Agnostic Yes